# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 4.2 Differentiability and minimisers of covariance loss function
Let $r\leq n$ and let $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ and $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ be as given in Lemma 3.6. Define
$\mathcal{P}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}^{-1}+\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{-1/2}w_{i},$ (20)
$\mathcal{C}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}-\sum_{i=1}^{r}-\lambda_{i}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}\otimes\mathcal{C}_{\textup{pr}}^{1/2}w_{i}.$ (21)
@@
Then $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are solutions to 4.4 and 4.3 respectively and $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are inverses of each other. For every $f\in\mathscr{F}$, the associated minimal loss is $\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}^{\textup{opt}}_{r})=\sum_{i>r}f(\lambda_{i})$. The solutions $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are unique if and only if either $\lambda_{r}=0$ or $\lambda_{r}<\lambda_{r+1}$.

### v3

@@ 4.4 Optimal low-rank posterior covariance approximations
Let $r\leq n$ and let $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ and $(w_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ be as given in Proposition 3.7. Define
$\mathcal{P}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}^{-1}+\sum_{i=1}^{r}\frac{-\lambda_{i}}{1+\lambda_{i}}(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})\otimes(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}),$ (21)
$\mathcal{C}^{\textup{opt}}_{r}$ $\coloneqq\mathcal{C}_{\textup{pr}}-\sum_{i=1}^{r}-\lambda_{i}(\mathcal{C}_{\textup{pr}}^{1/2}w_{i})\otimes(\mathcal{C}_{\textup{pr}}^{1/2}w_{i}).$ (22)
@@
Then $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are solutions to Problem 4.4 and Problem 4.3 respectively and $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are inverses of each other. For every $f\in\mathscr{F}$, the associated minimal loss is $\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}^{\textup{opt}}_{r})=\sum_{i>r}f(\lambda_{i})$. The solutions $\mathcal{P}^{\textup{opt}}_{r}$ and $\mathcal{C}^{\textup{opt}}_{r}$ are unique if and only if the following holds: $\lambda_{r+1}=0$ or $\lambda_{r}<\lambda_{r+1}$.
Remark 4.22 (Uniqueness condition).
Two remarks are in order when comparing the uniqueness characterisations of Theorem 4.21 and of its finite-dimensional analogue in [31, Theorem 2.3 and Corollary 3.1]. Firstly, the condition in Theorem 4.21 is not only sufficient but also necessary. Secondly, the sufficient condition of [31, Theorem 2.3 and Corollary 3.1] is that $(\frac{-\lambda_{1}}{1+\lambda_{1}},\ldots,\frac{-\lambda_{r}}{1+\lambda_{r}})$ are different, i.e. that $(\lambda_{1},\ldots,\lambda_{r})$ are different. From Theorem 4.21, we see that this condition should be interpreted as the condition that $(\lambda_{1},\ldots,\lambda_{r})$ are different from $\lambda_{r+1}$. Indeed, if $(\lambda_{1},\ldots,\lambda_{r})$ are different among each other but $\lambda_{r}=\lambda_{r+1}\not=0$, then replacing $(\lambda_{r},w_{r})$ by $(\lambda_{r+1},w_{r+1})$ in (22) and (21) gives a different solution to Problem 4.3 and Problem 4.4 respectively.